I am stuck on correct answer for:
what are all possible values of X such that X^2 = 5x

I am not using any specific book. I am just practicing..

Answer by anjulasahay(30)

I am stuck on correct answer for:
what are all possible values of X such that X^2 = 5x
ans : x^2-5x=0
x(x-5)=0
x = 0 or 5